# Comprehensive Evaluation of Prototype Neural Networks

### v1

@@ 8 Discussion
| CUB200 | CARS196 | NICO | AwA2 | |
ProtoPNet |
ProtoPool |
PIPNet |
In comparison with the results presented in the original papers, we find that the performance of our PIPNet models is approximately 8% lower than those reported by Nauta et al. [19]. We attribute this notable difference to the change in feature map resolution from in the original paper to used in our experiments to ensure a more equal comparison. The ProtoPNet and ProtoPool results exhibit a decrease of approximately 5% to 10% in the CUB200 and Cars196 datasets compared to the original paper. We attribute these differences primarily to the prototype channel reduction from 256 to 128, due to unstable performance results with the original channel configuration in our training setup and the imbalanced class distribution within our training set. Although class weights were employed in the classification loss to encourage equal learning across classes, cluster loss and separation loss were not similarly weighted. The original ProtoPNet and ProtoPool models were trained with an offline augmentation approach, which balanced the class distribution before training, making the use of weights unnecessary. Evaluating on a separate test set that was not utilized for selecting the best-performing model also generally results in a slight reduction in accuracy.
No comparable results exist for the NICO dataset. With performance approaching 90%, we conclude that prototype models exhibit general robustness to distribution shifts.
The performance of our AWA2 results ranges from 40% to 62%, with PIPNet demonstrating the highest performance and ProtoPNet Pruned the lowest. The relatively low performance across all models indicates that extracting understandable features is a challenging task for these models. Despite the frequently cited characterization of prototypes as general, interpretable features, the results from the AWA2 dataset demonstrate that there is substantial room for improvement in this context. We also acknowledge that performance can be further improved due to our inability to apply class weights in the classification loss despite a highly imbalanced label distribution, given that the applied multi-label margin loss does not accommodate this.
The compactness results from PIPNet indicate comparable reductions in global size while exhibiting high sparsity in the classification layer, supporting the findings reported in the original paper by Nauta et al. [19]. The results from ProtoPNet did not clearly verify the effects of pruning as discussed by Chen et al. [16], as seen by the marginal improvement of Background Overlap. Comparing the scores, it is evident that the non-negative classification layer of PIPNet is the only robust approach that guarantees positive reasoning among the models. Considering the sparsity and local size results, we conclude that PIPNet exhibits the highest interpretability in the classification layer.
The contrastivity results exhibit a similar pattern. PIPNet attains some of the highest results across the majority of metrics. Notably, the unique prototype locations without any overlap with other prototypes represent a significant improvement in interpretability over the other models. When we focus on the learned feature space, PIPNet shows well separated feature vectors underlying the effectiveness of the design. The other networks could not learn a meaningful latent space. This indicates that a well separated latent space is not robust to changing training setups.
The continuity evaluation demonstrates that PIPNet generally exhibits the most stable prototype locations and similarity scores, indicating that the network extracts stable high-level features rather than basic image characteristics. The other models frequently exhibit inferior results, suggesting that the contrastive learning strategy employed by PIPNet actively facilitates the extraction of high-level features.
In comparing the individual results of our output-completeness experiments, the PRP method exhibits a high degree of alignment with the subset of relevant pixels for a prototype within the PIPNet model. The other models exhibit inferior performance, suggesting that the PRP method is significantly influenced by the learned feature extraction. It can be argued that the learned feature extraction does not necessarily ensure the spatial alignment of the feature map with the input image. This suggests that visualizations can be enhanced by utilizing more sophisticated methods such as PRP rather than relying on the original upsampling method, which requires the spatial alignment between the feature map and the input image.
Wang et al. [58] report that the ProtoPNet model using the original upsampling visualization method achieves significantly higher measures on the CUB200 dataset. A visual comparison of the methods reveals that the upsampling method typically produces larger activation regions in the visualizations than the PRP method, thereby confirming these findings. The typical PRP visualization is illustrated in Figure 2, where the PRP saliency maps consist of multiple small regions.
The consistency results indicate that the prototypes are generally misaligned with specific object parts across the test set. Nauta et al. [19] introduced a purity metric to measure such alignment using the 10 most similar images to a prototype in the training set, where PIPNet’s prototypes achieve such an object-part alignment. Our consistency metric assigns an image to a prototype if it ranks among the top-5 most similar prototypes. Considering that the average local size of PIPNet is notably higher than 5, which implies that the prototypes in the top-5 exhibit reasonable similarity to the image. The results suggest that PIPNet’s object-part alignment decreases significantly in the absence of near-perfect similarity scores, which can only be observed in the 10 most similar images from the training set.

### v2

@@ 8 Discussion
| PIPNet | ProtoPNet | ProtoPNet Pruned | ProtoPool | ||
| (TP 11/17 TN 32/32) | (TP 17/17 TN 32/32) | (TP 17/17 TN 32/32) | (TP 17/17 TN 30/32) | ||
1. Proto. |
AwA2 |
2. Proto. |
| (correct pred.) | (correct pred.) | (correct pred.) | (correct pred.) | ||
1. Proto. |
Nico |
2. Proto. |
ProtoPNet primarily learned complex prototypes that captured low-level image characteristics rather than the intended semantic object parts, leading to an equally complex classification process. Pruning removed a small number of prototypes and slightly improved interpretability, but consistently deleting background prototypes could not be achieved.
Example visualizations on the fine-grained datasets (Figure 6) illustrate the challenges in interpreting the concepts learned by individual prototypes. While the learned concept can often be guessed when examining a single image, comparisons between the nearest feature visualizations from the training set and the sample image reveal that a prototype often encodes multiple concepts simultaneously. This issue is particularly pronounced in the other datasets (Figure 7). For example, in the AWA2 dataset, interpretation is especially vague, we can hardly guess prototype concepts such as “stripes” in zebras, while in the NICO dataset, sample features often span much larger regions than their nearest training features.
On CUB200 and Cars196, our accuracies are 5–10% below the original reports, most likely due to halving prototype channels (256→128) for more stability under our training setup. In addition, only the classification loss was class-weighted to counter imbalance; cluster/separation losses were not. Pruning produced only marginal gains in interpretability, falling short of the clearer pruning effects reported by Chen et al. [16]. Compared to Wang et al. [64], our PRP-based visualizations yield smaller, sparser activations than the original upsampling method, explaining their higher Object Overlap measures.
ProtoPool exhibited prototypes and classification processes of varying complexity, with clear differences between the two fine-grained datasets. Its performance appears to be strongly influenced by the complexity of shared object parts between classes. For example, in Cars196, shared car parts such as tires can be represented with low-level features like colour and shape, whereas in CUB200, the shared relations among bird species are more complex. Overall, ProtoPool tends to learn low-level image characteristics but emphasizes simpler, more interpretable classification processes than ProtoPNet, particularly in less complex classification tasks. We saw surprisingly good performance on the challenging multi-label dataset, suggesting that the shared prototype approach aligns well with such applications.
Figure 6 illustrates that prototypes in both fine-grained datasets often focus heavily on the background. Even in the better-performing Cars196 dataset, this issue is clearly visible. In contrast, visualizations on NICO and AWA2 (Figure 7) demonstrate improved object focus, comparable to ProtoPNet, which is consistent with ProtoPool’s stronger results on these datasets.
Similar to ProtoPNet, our ProtoPool scores on CUB200 and Cars196 are 5–10% below the original, that we explain by the same channel reduction and class-imbalance handling.
PIPNet achieved the best results in both prototype quality and classification simplicity, making it the most interpretable model. The combination of the softmax layer with the tanh loss proved highly effective in learning a minimal yet sufficient set of prototypes to achieve high accuracy. Furthermore, design choices in the classification layer, especially the hard positive constraint and the adapted output computation, successfully reduced the Local Size, yielding a more interpretable classification process compared to the other models.
However, visual explanations (Figures 6 and 7) reveal issues similar to those observed in ProtoPNet and ProtoPool. While prototype concepts can often be guessed in single-image visualizations, the nearest feature comparisons again reveal mismatches. Interestingly, PIPNet focuses more on background regions in the NICO and AWA2 datasets, showing the opposite trend of ProtoPool, despite both models employing shared prototypes. Relative to Nauta et al. [21], our PIPNet models underperform by 8%, which we primarily attribute to the lower feature map resolution used here ( vs. ) to ensure a uniform comparison across methods. Despite this gap, our compactness results mirror the original findings: comparable reductions in global size and high sparsity in the classification layer. PRP visualizations align best with PIPNet, though object-part consistency drops notably compared to the purity metric from the original study that focuses only on the top-10 nearest training images. On AWA2, PIPNet attains the highest accuracy. However, performance is likely suppressed by highly imbalanced labels and a multi-label margin loss that cannot incorporate class weights.
Overall, our evaluation focused on diverse model aspects relevant to interpretability, with mixed results across models and reappearing patterns in different evaluation techniques from individual architectures.